### v2

@@ 2.5 Receive signal model after CP removal
On the other hand, the transmission of the AU is adversely affected by intercarrier interference (ICI) due to the presence of Doppler shifts: in general, the channel matrix $\mathbf{H}_{\text{A},j}[n]$ cannot be diagonalized through DFT.
@@
2.6 Pilot allocation scheme
@@
$\mathcal{M}_{\text{TX}}\triangleq\{m_{\text{TX},0},m_{\text{TX},1},\ldots,m_{\text{TX},Q_{\text{TX}}-1}\}\subseteq\{0,1,\ldots,M-1\}$ (33)
@@
$\mathcal{N}_{\text{TX}}\triangleq\{n_{\text{TX},1},n_{\text{TX},2},\ldots,n_{\text{TX},N_{\text{TX, train}}}\}\subset\{0,1,\ldots,N_{0}-1\}\>.$ (34)
@@ 3.1 Blind estimation of Doppler and delay parameters
where the matrices in (35) are defined in (36)–(38) at the top of the next page, with ${\mathbf{\Delta}}\in\mathbb{C}^{M\times M}$ being a known diagonal matrix derived in the Appendix for the $\pi/2$-BPSK case and depending on the noncircularity property of the AU symbols.
${\mathbf{\Xi}}_{\text{A},j,\text{C}_{1},\text{C}_{2}}[-1]$ $\triangleq g_{\text{A},\text{C}_{1}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{1}})}\,g_{\text{A},\text{C}_{2}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{2}})}\,e^{\jmath\,2\pi\nu_{\text{A},\text{C}_{2}}}\,(\mathbf{D}_{\text{A},\text{C}_{1}}\,\mathbf{T}_{\text{A},\text{C}_{1}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{D}_{\text{A},\text{C}_{2}}\,\mathbf{T}_{\text{A},\text{C}_{2}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\in\mathbb{C}^{P\times P}$ (36)
@@
$\hskip 227.62204pt+(\mathbf{T}_{\text{A},\text{C}_{1}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A},\text{C}_{2}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}]\,\mathbf{D}_{\text{A},\text{C}_{2}}\in\mathbb{C}^{P\times P}$ (37)
${\mathbf{\Xi}}_{\text{A},j,\text{C}_{1},\text{C}_{2}}[1]$ $\triangleq g_{\text{A},\text{C}_{1}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{1}})}\,g_{\text{A},\text{C}_{2}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{2}})}\,e^{-\jmath\,2\pi\nu_{\text{A},\text{C}_{2}}}\,(\mathbf{D}_{\text{A},\text{C}_{1}}\,\mathbf{T}_{\text{A},\text{C}_{1}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{D}_{\text{A},\text{C}_{2}}\,\mathbf{T}_{\text{A},\text{C}_{2}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\in\mathbb{C}^{P\times P}$ (38)
@@ 3.1.1 Blind estimation of the Doppler shifts
In practice, the estimates of the Doppler shifts can be directly obtained from data by finding the peaks of the function $\widehat{J}(\alpha)$ obtained from (44) by replacing (40) with (43). Consistency and asymptotic normality of $\widehat{J}(\alpha)$ can be proven by paralleling the derivations detailed in [53].
@@ 3.1.2 Blind estimation of the time delays
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}\,(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}$ (47)
@@
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}[(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}$
$\hskip 56.9055pt+(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}]$ (48)
@@
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}\,(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\>.$ (49)
@@
In practice, an estimate $\widehat{\tau}_{\text{A,C}}$ of the delay $\tau_{\text{A,C}}$ is derived by searching for the peak of the finite-sample version ^$\widehat{\mathcal{I}}_{\text{A,C}}(\beta)$ of the cost function $\mathcal{I}_{\text{A,C}}(\beta)$ defined in (60), where ^$\widehat{\mathcal{I}}_{\text{A,C}}(\beta)$ is obtained by replacing ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\nu_{\text{A,C}}}[r]$ in (47)–(49), for $r\in\{-1,0,1\}$, with their corresponding estimates $\widehat{{\mathbf{R}}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\widehat{\nu}_{\text{A,C}}}[r]$ [see (43)], evaluated at the estimated cycle frequency $2\,\widehat{\nu}_{\text{A,C}}$ acquired through the algorithm in Subsection II-A-1. The performance of the proposed estimator for time delays is assessed by simulations in Section 5.
@@ 3.2 Pilot-aided estimation of complex gain and AoA parameters
After acquiring the Doppler and delay parameters blindly, we propose to estimate the complex gain and AoA unknowns of the AU, by capitalizing on its pilots transmitted in the symbol training intervals $\mathcal{N}_{\text{A}}=\{n_{\text{A},1},n_{\text{A},2},\ldots,n_{\text{A},N_{\text{A,train}}}\}$. For the present, Doppler and delay parameters are assumed to be perfectly known. The starting point of the pilot-aided estimation step is the signal model (22), which collects the signal received by the $j$-th antenna after CP removal and DFT.
@@
${\mathcal{L}}_{A}(\rho_{\text{L}},\rho_{\text{N}},\varrho_{\text{L}},\varrho_{\text{N}})\triangleq\sum_{j=1}^{J}\left\|{\mathbf{Y}}_{\text{A},j}-\sum_{\text{C}\in\{\text{L, N}\}}\rho_{\text{C}}\,\,\varrho_{\text{C}}^{j-1}\,\mathbf{P}_{\text{A,C}}\right\|^{2}\>.$ (66)
@@
In practice, the previously obtained estimates of the Doppler and delay parameters (see Subsection III-A-1 and III-A-2) have to be used for LS acquisition of the complex gain and AoA unknowns of the AU. Specifically, the LS problem to be solved comes from replacing $\mathbf{P}_{\text{A,C}}$ in (66) with its corresponding estimate $\widehat{\mathbf{P}}_{\text{A,C}}\triangleq(\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},1}],\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},2}],\ldots,\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},N_{\text{A,train}}}])$, where $\widehat{\mathbf{p}}_{\text{A,C}}[n]\triangleq{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{R}}_{\text{cp}}\,\widehat{\mathbf{D}}_{\text{A,C}}\,\widehat{\mathbf{T}}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}}\,{\mathbf{s}}_{\text{A}}[n]\,e^{\jmath\,2\pi\widehat{\nu}_{\text{A,C}}n}$, for any $n\in\mathcal{N}_{\text{A}}$, and $\widehat{\mathbf{T}}_{\text{A,C}}^{(0)}$ is built from (45) by using $\widehat{\tau}_{\text{A,C}}$ in lieu of $\tau_{\text{A,C}}$. The performance of this LS estimator is assessed by simulations in Section 5.
@@ 5 Numerical results
where $\Delta_{\text{max}}=3\,T_{\text{c}}$, the slope-time is $\tau_{\text{slop}}=2\,T_{\text{c}}$, and $u$ is a random variable uniformly distributed into $[0,1]$. We assume that the constant radial speed of the AU is $v=6$ m/s, $f_{\text{max}}=720$ Hz, and the AoD $\vartheta_{\text{A,L}}$ and $\vartheta_{\text{A,N}}$ are independent random variables uniformly distributed into $[0,\pi]$. The power relationship between the AU and TU at the BS is measured by the aerial-to-terrestrial ratio (ATR) defined as
@@
In each Monte Carlo run, a new set of symbols, noise and channel parameters for both the AU and TU are randomly generated. Regarding the estimation of the channel parameters of the AU (see Section 3), we calculate the arithmetic mean of the MSEs of the Doppler shifts $f_{\text{A,L}}$ and $f_{\text{A,N}}$ (normalized by $f_{\text{max}}^{2}$), the time delays $\tau_{\text{A,L}}$ and $\tau_{\text{A,N}}$ (normalized by $\Delta^{2}_{\text{max}}$), the channel gains $g_{\text{A,L}}$ and $g_{\text{A,N}}$ (normalized by $\sigma^{2}_{\text{A,L}}$), and the direction cosines $\sin(\theta_{\text{A,L}})$ and $\sin(\theta_{\text{A,N}})$. With reference to the estimation of the TU channel (see Section 4), we report the estimator output (97) (normalized by $\sum_{k=1}^{K_{\text{T}}}\sigma_{g_{\text{T},k}}^{2}$).
@@
Figure 7: MSE of TU channel versus SNR and ATR values (dotted line for the LS estimator and solid line for the BWLU estimator).
@@
Fig. 2 reports the MSE of the Doppler shifts as a function of the length $N_{0}$ of the observation interval (in OFDM symbols). As expected, the estimation accuracy rapidly improves as $N_{0}$ increases, thus demonstrating that the proposed blind estimator of the Doppler shifts is asymptotically consistent. Remarkably, Fig. 3 shows that the MSE of the Doppler shifts is independent of both the SNR and ATR, achieving the value $-62.57$ dB approximately. Such an invariance with respect to both the SNR and ATR comes from the fact that the blind estimator proposed in Subsection III-A-1 relies on the second-order conjugate cyclic statistics of the received data, for which the contribution of the circular wide-sense stationary TU signal and noise is negligible for sufficiently large values of $N_{0}$.
@@
We remember that, as the Doppler shifts, the time delays are blindly estimated by using the conjugate second-order statistics of the ACS random process (17). In this case, a weak dependence of the MSE on the SNR and ATR values can be observed in Fig. 4, which is due to the fact that the estimates of the time delays in Subsection III-A-2 also feel the aftereffects of the estimation error of the cycle frequency acquired previously. However, the variability of the MSE of the time delays is confined in a range of about $1$ dB as a function of the SNR and the MSE curves translate down of nearly $1$ dB further to an ATR reduction of $3$ dB.
The MSE of the channel amplitudes in Fig. 5 also exhibits a weak dependence on the SNR and ATR values. However, the additional exploitation of the pilot symbols allows to achieve MSE values smaller than $-10$ dB for $\text{SNR}>4$ dB, even when the ATR is as low as $-3$ dB. In contrast, the MSE of the AoAs in Fig. 6 is actually independent of the ATR, thus further corroborating the robustness of the proposed pilot-based estimators developed in Subsection III-B against the interference generated by the TU.
@@
Fig. 7 depicts the MSE performance of the TU channel estimator proposed in Section 4. Besides the MSE of the BWLU estimator (98) (solid line), we also report the MSE of the simpler LS estimator (87) (dotted line). As predicted by our analysis, the BWLU estimator largely outperforms the LS one since it additionally exploits the CSI knowledge of the AU. This demonstrates the importance of firstly estimate the channel parameters of the AU and, then, using the obtained estimates to acquire the CSI of the TU.
6 Conclusions

### v4

@@ 2.5 Receive signal model after CP removal
On the other hand, the transmission of the AU is adversely affected by ICI due to the presence of Doppler shifts: in general, the channel matrix $\mathbf{H}_{\text{A},j}[n]$ cannot be diagonalized through DFT.
@@
2.6 Nonorthogonal pilot allocation scheme
@@
$\mathcal{M}_{\text{TX}}\triangleq\{m_{\text{TX},0},m_{\text{TX},1},\ldots,m_{\text{TX},Q_{\text{TX}}-1}\}\\ \subseteq\{0,1,\ldots,M-1\}$ (33)
@@
$\mathcal{N}_{\text{TX}}\triangleq\{n_{\text{TX},1},n_{\text{TX},2},\ldots,n_{\text{TX},N_{\text{TX, train}}}\}\\ \subset\{0,1,\ldots,N_{0}-1\}\>.$ (34)
@@ 3 Channel estimation for the aerial user
The proposed estimators of the AU channel parameters are summarized in Fig. 2.
Figure 2: A summary of the proposed channel estimators for both the AU and TU (the BWLU estimator for the TU channel is depicted only).
@@ 3.1 Blind estimation of Doppler and delay parameters
where the $P\times P$ matrices in (35) are defined in (36)–(38) at the top of the next page, with ${\mathbf{\Delta}}\in\mathbb{C}^{M\times M}$ being a known diagonal matrix derived in the Appendix for the $\pi/2$-BPSK case and depending on the noncircularity property of the AU symbols.
${\mathbf{\Xi}}_{\text{A},j,\text{C}_{1},\text{C}_{2}}[-1]$ $\triangleq g_{\text{A},\text{C}_{1}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{1}})}\,g_{\text{A},\text{C}_{2}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{2}})}\,e^{\jmath\,2\pi\nu_{\text{A},\text{C}_{2}}}\,(\mathbf{D}_{\text{A},\text{C}_{1}}\,\mathbf{T}_{\text{A},\text{C}_{1}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{D}_{\text{A},\text{C}_{2}}\,\mathbf{T}_{\text{A},\text{C}_{2}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}$ (36)
@@
$\hskip 227.62204pt+(\mathbf{T}_{\text{A},\text{C}_{1}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A},\text{C}_{2}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}]\,\mathbf{D}_{\text{A},\text{C}_{2}}$ (37)
${\mathbf{\Xi}}_{\text{A},j,\text{C}_{1},\text{C}_{2}}[1]$ $\triangleq g_{\text{A},\text{C}_{1}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{1}})}\,g_{\text{A},\text{C}_{2}}\,e^{\jmath\pi(j-1)\sin(\theta_{\text{A},\text{C}_{2}})}\,e^{-\jmath\,2\pi\nu_{\text{A},\text{C}_{2}}}\,(\mathbf{D}_{\text{A},\text{C}_{1}}\,\mathbf{T}_{\text{A},\text{C}_{1}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{D}_{\text{A},\text{C}_{2}}\,\mathbf{T}_{\text{A},\text{C}_{2}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}$ (38)
@@ 3.1.1 Blind estimation of the Doppler shifts
In practice, the estimates of the Doppler shifts can be obtained from data by finding the peaks of the function $\widehat{J}(\alpha)$ obtained from (44) by replacing (40) with (43).
@@ 3.1.2 Blind estimation of the time delays
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}$
$\hskip 56.9055pt\cdot(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}$ (47)
@@
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}$
$\cdot\left[(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\right.$
$\left.\hskip 42.67912pt+(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\right]$ (48)
@@
$=g_{\text{A,C}}^{2}\,e^{\jmath 2\pi(j-1)\sin(\theta_{\text{A,C}})}$
$\hskip 56.9055pt\cdot(\mathbf{T}_{\text{A,C}}^{(1)}\,{\mathbf{\Omega}}_{\text{A}})\,{\mathbf{\Delta}}\,(\mathbf{T}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}})^{\text{T}}\>.$ (49)
@@
In practice, an estimate $\widehat{\tau}_{\text{A,C}}$ of the delay $\tau_{\text{A,C}}$ is derived by searching for the peak of the finite-sample version ^$\widehat{\mathcal{I}}_{\text{A,C}}(\beta)$ of the cost function $\mathcal{I}_{\text{A,C}}(\beta)$ defined in (60), where ^$\widehat{\mathcal{I}}_{\text{A,C}}(\beta)$ is obtained by replacing ${\mathbf{R}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\nu_{\text{A,C}}}[r]$ in (47)–(49), for $r\in\{-1,0,1\}$, with their corresponding estimates $\widehat{{\mathbf{R}}}_{\overline{{\mathbf{y}}}_{j}\overline{{\mathbf{y}}}_{j}^{*}}^{2\,\widehat{\nu}_{\text{A,C}}}[r]$ [see (43)], evaluated at the estimated cycle frequency $2\,\widehat{\nu}_{\text{A,C}}$ acquired through the algorithm in Subsection II-A-1.
@@ 3.2 Pilot-aided estimation of complex gain and AoA parameters
After acquiring the Doppler and delay parameters blindly, we propose to estimate the complex gain and AoA unknowns of the AU, by capitalizing on its pilots transmitted in the training intervals $\mathcal{N}_{\text{A}}=\{n_{\text{A},1},n_{\text{A},2},\ldots,n_{\text{A},N_{\text{A,train}}}\}$. For the present, Doppler and delay parameters are assumed to be perfectly known. The starting point of the pilot-aided estimation step is (22), which collects the signal received by the $j$-th antenna after CP removal and DFT.
@@
${\mathcal{L}}_{A}(\rho_{\text{L}},\rho_{\text{N}},\varrho_{\text{L}},\varrho_{\text{N}})\triangleq\sum_{j=1}^{J}\Big{\|}{\mathbf{Y}}_{\text{A},j}-\sum_{\text{C}\in\{\text{L, N}\}}\rho_{\text{C}}\,\,\varrho_{\text{C}}^{j-1}\,\mathbf{P}_{\text{A,C}}\Big{\|}^{2}\>.$ (66)
@@
In practice, the previously obtained estimates of the Doppler and delay parameters (see Subsection III-A-1 and III-A-2) have to be used for LS acquisition of the complex gain and AoA unknowns of the AU. Specifically, the LS problem to be solved comes from replacing $\mathbf{P}_{\text{A,C}}$ in (66) with its corresponding estimate $\widehat{\mathbf{P}}_{\text{A,C}}\triangleq(\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},1}],\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},2}],\ldots,\widehat{\mathbf{p}}_{\text{A,C}}[n_{\text{A},N_{\text{A,train}}}])$, where $\widehat{\mathbf{p}}_{\text{A,C}}[n]\triangleq{\mathbf{W}}_{M}^{\text{H}}\,{\mathbf{R}}_{\text{cp}}\,\widehat{\mathbf{D}}_{\text{A,C}}\,\widehat{\mathbf{T}}_{\text{A,C}}^{(0)}\,{\mathbf{\Omega}}_{\text{A}}\,{\mathbf{s}}_{\text{A}}[n]\,e^{\jmath\,2\pi\widehat{\nu}_{\text{A,C}}n}$, for any $n\in\mathcal{N}_{\text{A}}$, and $\widehat{\mathbf{T}}_{\text{A,C}}^{(0)}$ is built from (45) by using $\widehat{\tau}_{\text{A,C}}$ in lieu of $\tau_{\text{A,C}}$.
3.3 Computational issues
With reference to Fig. 2, the computational complexity of the overall estimation algorithm for the AU is mainly dominated by the first block devoted to the Doppler shifts acquisition, i.e., the CCCM estimation. In particular, the CCCM estimation entails $\mathcal{O}(N_{0}^{2}\,P^{2}\,J)$ floating point operations (flops) and, thus, the complexity burden quadratically grows with the length $N_{0}$ of the observation interval.
3.4 NOMA is used by TUs in the uplink
So far, we have assumed that the TUs transmit to the BS using the OMA scheme. We now discuss how the proposed channel estimation algorithm for the AU will work if nonorthogonal RBs are assigned to TUs in the uplink.
In this scenario, a single AU may send noncircular symbols using the same RBs shared by a group of static TUs that concurrently transmit to the BS by employing a power-domain NOMA scheme with circular modulation formats. In such a case, the proposed blind estimation procedure of Doppler and delay parameters is unaffected, since the terrestrial transmissions do not contribute to the CCCM (41) due to the modulation and Doppler diversities. On the other hand, the uplink signals of the TUs appear in the disturbance term ${\mathbf{d}}_{\text{A},j}[n]$ in (61), thus impacting on the pilot-aided estimation of complex gain and AoA parameters. However, existing schemes in the power-domain NOMA literature (e.g., [45, 66]) creates multiple groups, each with one “weak” TU at the cell edge and one “strong” TU at the cell center. For such a NOMA scheme, the power of ${\mathbf{d}}_{\text{A},j}[n]$ will be mainly dictated by the strong TU and, hence, compared with the single TU case considered in this paper, the performance degradation of the proposed LS estimator (65) will be negligible.
@@ 4 Channel estimation for the terrestrial user
4.1 Computational issues
With reference to Fig. 2, the computational complexity of the estimation algorithm for the TU is mainly dominated by the second block of the channel acquisition for the TU, i.e., the BWLU estimation. In particular, the BWLU estimation relies on the inversion of the matrix ${\mathbf{R}}_{\widetilde{{\mathbf{d}}}_{\text{T}}\widetilde{{\mathbf{d}}}_{\text{T}}}$ which entails $\mathcal{O}(J^{3}\,Q^{3}_{\text{T}}\,N^{3}_{\text{T,train}})$ flops if one resorts to batch algorithms. Such a matrix inversion can also be implemented by means of a simple and effective recursion, similar to the well-known recursive least square algorithm, with a complexity per iteration of order only $\mathcal{O}(J^{2}\,Q^{2}_{\text{T}}\,N^{2}_{\text{T,train}})$ flops. Anyway, the complexity burden is independent of the length $N_{0}$ of the observation interval.
4.2 NOMA is used by TUs in the uplink
As in Subsection III-3.4, we discuss herein how the proposed channel estimation algorithms for the TU will be employed when the BS assigns to a single AU the same RBs shared by a group of power-domain NOMA TUs.
If the BS uses orthogonal pilot sequences to distinguish different power-domain NOMA TUs [67], then the received signal from a given TU can be obtained by performing the correlation of the received data with the pilot sequence of the desired TU. After performing this preprocessing for each user, the estimators (87) or (96) will be used with minor modifications to individually estimate the channel of each TU. On the other hand, if the BS allocates the same pilot to the power-domain NOMA TUs, the estimators (87) or (96) will estimate a linear combination of the channels from all NOMA TUs. Such an estimate will still provide a useful description of the combined channel, which could be exploited by the BS to beamform a combination of the data symbols intended for the NOMA TUs [66].
@@ 5 Numerical results
where $\Delta_{\text{max}}=3\,T_{\text{c}}$, the slope-time is $\tau_{\text{slop}}=2\,T_{\text{c}}$, and $u$ is a random variable uniformly distributed into $[0,1]$. Unless otherwise specified, we assume that the constant radial speed of the AU is $v=8$ m/s, $f_{\text{max}}=720$ Hz, and the AoD $\vartheta_{\text{A,L}}$ and $\vartheta_{\text{A,N}}$ are independent random variables uniformly distributed into $[0,\pi]$. The power relationship between the AU and TU at the BS is measured by the aerial-to-terrestrial ratio (ATR) defined as
@@
In each one of the $400$ Monte Carlo runs, a new set of symbols, noise and channel parameters for both the AU and TU are randomly generated. Regarding the estimation of the channel parameters of the AU (see Section 3), we calculate the arithmetic mean of the MSEs of the Doppler shifts $f_{\text{A,L}}$ and $f_{\text{A,N}}$ (normalized by $f_{\text{max}}^{2}$), the time delays $\tau_{\text{A,L}}$ and $\tau_{\text{A,N}}$ (normalized by $\Delta^{2}_{\text{max}}$), the channel gains $g_{\text{A,L}}$ and $g_{\text{A,N}}$ (normalized by $\sigma^{2}_{\text{A,L}}$), and the direction cosines $\sin(\theta_{\text{A,L}})$ and $\sin(\theta_{\text{A,N}})$. With reference to the estimation of the TU channel (see Section 4), we report the estimator output (95) (normalized by $\sum_{k=1}^{K_{\text{T}}}\sigma_{g_{\text{T},k}}^{2}$). As baseline channel estimation schemes, we report the OMA-based counterparts of the proposed estimators, which are implemented by orthogonalizing pilot symbols in the frequency domain, i.e., we set $\mathcal{M}_{\text{A}}\equiv\{0,1,\ldots,7\}$ and $\mathcal{M}_{\text{T}}\equiv\{8,9,\ldots,15\}$ in each training symbol interval.
@@
Figure 8: MSE of TU channel versus SNR and ATR values (dashed line for the LS-NOMA estimator, solid line for the BWLU estimator, and dotted line for the LS-OMA estimator).
@@
Fig. 3 reports the MSE of the Doppler shifts as a function of the length $N_{0}$ of the observation interval (in OFDM symbols). As expected, the estimation accuracy rapidly improves as $N_{0}$ increases, thus demonstrating that the proposed blind estimator of the Doppler shifts is asymptotically consistent. Remarkably, Fig. 4 shows that the MSE of the Doppler shifts is independent of both the SNR and ATR, achieving the value $-62.57$ dB approximately. Such an invariance with respect to both the SNR and ATR comes from the fact that the blind estimator proposed in Subsection III-A-1 relies on the second-order conjugate cyclic statistics of the received data, for which the contribution of the circular wide-sense stationary TU signal and noise is negligible for sufficiently large (but finite) values of $N_{0}$.
@@
We remember that, as the Doppler shifts, the time delays are blindly estimated by using the conjugate second-order statistics of the ACS random process (17). In this case, a weak dependence of the MSE on the SNR and ATR values can be observed in Fig. 5, which is due to the fact that the estimates of the time delays in Subsection III-A-2 also feel the aftereffects of the estimation error of the cycle frequency acquired previously. Estimation errors due to finite sample-size effects (i.e., when the conjugate second-order statistics are estimated from a finite number of samples) depend on both the SNR and ATR in a nonlinear manner [68], thus contributing to the fluctuations of the curves in Fig. 5. However, the variability of the MSE of the time delays is confined in a range of about $1$ dB as a function of the SNR and the MSE curves translate down of nearly $1$ dB further to an ATR reduction of $3$ dB. It is worth noticing that the performance of the proposed NOMA-based estimators in Figs. 4 and 5 is almost indistinguishable from that of the corresponding OMA-based counterparts, thus corroborating the high robustness of the proposed Doppler shifts and time delays estimation approaches against the interference caused by the TU.
The MSE of the channel amplitudes in Fig. 6 also exhibits a weak dependence on the SNR and ATR values. However, the additional exploitation of the pilot symbols allows to achieve MSE values smaller than $-10$ dB for $\text{SNR}>2$ dB, even when the ATR is as low as $-3$ dB. In contrast, the MSE of the AoAs in Fig. 7 is actually independent of the ATR, thus further corroborating the robustness of the proposed pilot-based estimators developed in Subsection III-B against the interference generated by the TU. With respect to Figs. 4 and 5, fluctuations in the MSE curves of channel amplitudes in Fig. 6 and AoAs in Fig. 7 become even more evident since, although the corresponding estimation procedures are based on a LS approach that does not involve estimation of statistics, these estimators suffer of the error propagation phenomenon that is inherent of the proposed multi-stage estimation approach (see Fig. 2), according to which the existence of estimation error in a given stage will deteriorate the estimation accuracy in the next stage. Such an error propagation prevents the LS estimates of channel amplitudes and AoAs from monotonically decreasing for increasing values of SNR and/or ATR. The proposed NOMA-based estimators and their corresponding OMA-based counterparts in Figs. 6 and 7 exhibit slight performance differences. In particular, the performance of OMA-based estimators gets slightly worse than that of the corresponding NOMA-based counterparts in the case of amplitude AU estimation. This behavior is due to the fact that amplitude estimation procedures are typically very sensitive to the number of available pilot symbols. Indeed, as a consequence of the orthogonalization of pilot symbols in the frequency domain, OMA-based estimators count on a reduced number of pilot symbols, compared to NOMA-based estimation algorithms.
@@
Fig. 8 depicts the MSE performance of the TU channel estimator proposed in Section 4. Besides the MSE of the BWLU estimator (96) (solid line), we also report the MSE of the simpler LS-NOMA estimator (87) (dashed line). As predicted by our analysis, the BWLU estimator largely outperforms the LS-NOMA one since it additionally exploits the CSI knowledge of the AU. This demonstrates the importance of firstly estimate the channel parameters of the AU and, then, using the obtained estimates to acquire the CSI of the TU. As a matter of fact, the performance curves of the OMA-based LS estimator (dotted line) are monotonically decreasing functions of the SNR, whereas the MSE performance of the corresponding NOMA-based counterparts exhibits a floor in the high SNR regime, which is due to the residual interference from the AU transmission.
Table 1: MSEs of AU and TU channel parameters for different values of the radial speed of the AU [dB] .
$v$ [m/s] Doppler Shifts AU Time Delays AU Amplitude AU AoA AU Channel TU
Proposed OMA Proposed OMA Proposed OMA Proposed OMA Proposed LS-NOMA LS-OMA
2 -62.8159 -62.8159 -11.2695 -11.2695 -16.2791 -15.7644 -2.7585 -3.9616 -37.4665 -24.9258 -34.1684
4 -62.6253 -62.6253 -10.8137 -10.8137 -15.9419 -15.7462 -3.3521 -4.5738 -37.1741 -24.5377 -34.2129
8 -62.5776 -62.5776 -10.3907 -10.3907 -14.5273 -14.1755 -4.6412 -5.1795 -36.5062 -24.5013 -34.3215
16 -62.5230 -62.5230 -9.7083 -9.7083 -12.8208 -12.5091 -6.7672 -6.0224 -35.2183 -24.3936 -34.8025
Finally, Tab. 1 reports the MSE performance of all the estimation algorithms under comparison for different value of the radial speed $v$ of the AU. It can be inferred that the radial speed of the AU has a negligible effect on the estimates, except for the estimated values of the AoA of the AU, whose accuracy improves for increasing value of $v$.
6 Conclusions and directions for future work